# Technological Complexity Based on Japanese Patent Data

### v3

@@ Abstract
As international competition intensifies in technologies, nations need to identify key technologies to foster innovation. However, the identification is challenging due to the independent and inherently complex nature of technologies. Traditionally, regional analyses of technological portfolios have been limited to binary evaluations, indicating merely whether a region specializes in a technology, or relying on the average Technological Complexity Index (TCI) of the specialized technologies. This study proposes that evaluating TCI at the corporate level could provide finer granularity and more detailed insights. To address the underutilization of corporate-level TCI assessments in Japan, this study applies the Technological Complexity Index using carefully processed patent data spanning fiscal years 1981 to 2010. Specifically, we analyze a bipartite network composed of 1,938 corporations connected to technological fields categorized into either 35 or 124 classifications. Our findings quantitatively characterize the ubiquity and sophistication of each technological field, reveal detailed technological trends reflecting broader societal contexts, and demonstrate methodological stability even when employing finer technological classifications. Additionally, our corporate-level approach allows consistent comparisons across different regions and technological fields, clarifying regional advantages in specific technologies. This refined analytical framework offers policymakers and researchers robust, targeted insights, thereby significantly contributing to innovation strategy formulation in Japan.
@@
Keyword1, Keyword2, Keyword3
@@ Introduction
In the twenty-first century, intensifying international competition has made technological capability a central element of national competitiveness. Nations seek to foster environments that support the development of sophisticated technologies, as these capabilities are fundamental to sustainable economic growth. [1, 2, 3] The ecosystems that produce such capabilities are, however, inherently complex, involving interactions among countries, corporations, and technologies [4]. This complexity makes it difficult to accurately identify the key technologies and actors within the system. Understanding this structure is crucial for developing effective science and technology strategies to strengthen a nation’s competitive advantage.
@@
Historically, the assessment of technological capabilities has predominantly relied on quantitative metrics derived from bibliographic data of patents and publications, ranging from simple patent counts to more complex citation-based indicators, including Technology Relevance (TR) [5, 6]. While such indicators offer certain insights by considering citation practices, patent age, and technological fields, they suffer from significant shortcomings. Their inherent dependence on the long-term accumulation of citations makes them susceptible to biases from citation behaviors, patent examiners, and strategic patenting. Crucially, these metrics fail to capture the core aspect of technological complexity, namely the intricate interplay among corporations and technologies. Consequently, approaches that rely solely on these data provide a limited perspective and inadequately reflect the qualitative dimensions of technological capability.
@@
The HH algorithm provides a powerful solution by considering both the diversity of a region and the ubiquity of its products or technologies [7]. This dual perspective allows the methodology to distinguish between components and quantify their complexity, revealing detailed economic structures not captured by traditional metrics. The utility of this foundational framework has led to its extension into several distinct streams of research[8, 9, 10, 11, 12]. One major development is the introduction of complementary metrics such as fitness-based approaches [13, 14, 15]. Another significant stream, which is the focus of this study, involves the adaptation of the algorithm to patent data, pioneered by Balland and Rigby to measure technological complexity [16]. Building on this work, recent studies have further extended the patent-based analysis to contexts such as Asian economies [17, 18]. Despite this wide range of applications, the focus has predominantly remained at a macro level, such as on cities or regions [8]. Consequently, even within the stream of patent analysis, comprehensive validation at the corporation level is still limited.
@@
To address this gap, our study applies the HH algorithm to an extensive set of Japanese patent records to quantify technological complexity at the corporation level. By constructing bipartite networks linking corporations to technology classes, we develop a detailed measure of technological sophistication. Our analysis investigates the relationships between technological complexity, ubiquity, and the average corporate diversity of patent technologies, providing new insights into the nature of advanced capabilities. The model developed at the corporation level provides a more robust and consistent assessment of technology than regional-level models, particularly for finely classified technological fields. Importantly, this study shows that a corporation-level approach enables a direct comparison of technological characteristics across different regions. By focusing on corporations within specific locales, we can identify distinct industrial specializations and reveal which technologies are complex within each region’s unique context. The framework and insights from this study provide a powerful tool for creating tailored industrial policies and more accurately identifying the key technologies vital for future growth.
@@
The rest of this paper is organized as follows. In Introduction section, we present the ubiquity and complexity of technological fields presented by the patent data, followed by a detailed analysis of technological trends. In the Finer Technological Trends and Consistency section, we provide a critical analysis of our findings and discuss the implications of our results. Finally, in the Discussion section, we detail the data processing and methods used in our analysis.
@@ Results
In this study, we use the HH algorithm to quantify the structural characteristics of technologies based on patent publications from 1,938 corporations filed between fiscal years 1981 and 2010 (see Discussion section). First, we define ubiquity (or degree centrality) $K_{T,0}$ as the number of corporations connected to a technology which lower value indicates that the technology is more rare. Similarly, the average diversity $K_{T,1}$, which measures the variety of corporations connected to a technology, implies that a higher value reflects diversified publication, whereas a lower value suggests specialization. Generally, the average of each indicator is used as a threshold to distinguish high from low values [19, 16, 20].
@@
To systematically evaluate technological characteristics, we map the relationships among ubiquity $K_{T,0}$, average diversity $K_{T,1}$, and TCI across patent technologies. (Fig. 1) Technologies are classified according to the Schmoch system [21], as illustrated in Panel A of Fig. 1. Values for ubiquity and average diversity are normalized by their respective means, while TCI is segmented by the zero threshold.
@@
Detailed analysis using International Patent Classification (IPC) classes further supports these observations (Fig. A1). Within the five major sectors defined by Schmoch [21], the relationship between ubiquity and average diversity is predominantly weak or moderate. Specifically, Electrical Engineering displays a strong positive correlation between ubiquity and TCI (Pearson correlation coefficient $r=0.738$), highlighting the coexistence of corporations specializing in broadly applicable technologies and those focusing on niche, specialized innovations.
@@
In contrast to Electrical engineering technologies, Chemistry and Pharmaceuticals technologies mostly exhibit high TCI and the high average diversity (Panel D in Fig. 1, Pearson correlation coeffecience $r=0.316$). These sectors are characterized by rarity and high sophistication and are typically developed by more diversified corporations. Moreover, corporations in these domains are connected not only to fields with low ubiquity but also to those that are more common (e.g., Biotechnology, Analysis of Biological Materials, and Measurement as shown in the right quadrant of Panel C in Fig. 1). A detailed examination reveals a slight negative correlation between ubiquity and average diversity (Fig. A1), suggesting that in sectors with high technological complexity, corporations with more comprehensive technological portfolios tend to produce rare technologies. Furthermore, the IPC classes corresponding to Biotechnology and Measurement, show high TCI values in fields related to the production of sugar and yeast (C13, A21, C12), drug discovery (A61), and the preservation and management of these technologies (C07, A23). These findings are consistent with the socio-technical context of the period, particularly reflecting the emergence of drug discovery technologies in the 1980s and 1990s [22, 23]. Moreover, the significant roles of food chemistry and biotechnology during the 1990s and 2000s underscore their contribution to the development of additives that enabled safer food preservation and processing not only in Japan but globally [24, 25].
Figure 1: (A) The legend of the five classifications defined by Schmoch [21], along with an explanation of the technology codes used in (B), (C), and (D). (B) Pearson correlation between $K_{T,0}$ and $K_{T,1}$ ($r=0.064$). (C) Pearson correlation between $K_{T,0}$ and TCI ($r=0.594$). (D) Pearson correlation between TCI and $K_{T,1}$ ($r=0.316$). The black lines indicate the mean values of $K_{T,0}$ and $K_{T,1}$, and $\text{TCI}=0$ is also shown with a black line. Each dot is color-coded according to (A).
Finer Technological Trends and Consistency
The traditional regional approach measures technological complexity as a relative index reflecting regional technological capabilities, shaped by the economic advantages arising when corporations cluster within particular countries or cities, known as agglomeration economies [26]. Consequently, this method inherently captures spatial inequalities and is strongly influenced by urbanization levels in Japan. In terms of evaluating a region’s technological portfolio, this approach typically provides limited insights, either indicating binary specialization in technologies (0 or 1) or offering only the average TCI of the specialized technologies. Furthermore, finer classifications, such as IPC classes, tend to produce less consistent complexity indices compared to broader classifications like the Schmoch system [27]. Such inconsistencies appear in both regional and technological complexity evaluations due to the symmetric nature of the bipartite network structure underlying these calculations.
In this study, we propose examining technological complexity at the corporate level to overcome these limitations and achieve a more detailed understanding of regional technological portfolios. Specifically, by calculating TCI at the corporate level, we explore whether it is possible to capture more granular insights into technological capabilities within regions. By comparing TCI results derived from regional and corporate approaches, we assess how classification granularity influences the evaluation of inequality and stability of technological complexity across regions and corporations (Panel A in Fig. 2). In this study, we address these inconsistencies by comparing TCI results obtained from regional and corporate approaches, analyzing how classification granularity affects inequality and stability of complexity across regions and corporations (Panel A in Fig. 2).
The scaled TCI of the Schmoch system shows that in the regional approach 26 out of 35 classes fall in the range of 75 to 100. In the corporate approach, only 15 classes lie in that range, whereas the other classes display greater differences among technologies in the regional approach. This difference in stems from the corporate approach being able to characterize each technology in a more detailed manner. The HH algorithm defines ubiquity as the degree of a technology node and this measure depends on the number of paired nodes. The regional approach comprised roughly a few hundred region nodes, such as countries or EU cities[8, 28]. The corporate approach has corporation nodes numbering from one thousand to several ten thousand. The iterative calculation of TCI can therefore reflect a wider range of ubiquity.
The difference between the approaches also appears when using a finer classification. When the classification becomes more detailed and the number of technology nodes increases, the regional approach cannot define ubiquity sufficiently. In the case of Civil engineering the variance of TCI values for IPC classes corresponding to the same Schmoch class is larger than that for the corporate approach. The larger difference between the TCI values of the coarse classification and those of the finer classification indicates a greater inconsistency in the evaluation for the regional approach. A Wilcoxon signed rank test shows that the distribution of residuals between the TCI values of the Schmoch and IPC classes is significantly different (p value = 7.23e-14) and that the residuals for the corporate approach are smaller than those for the regional approach (Panel B in Fig. 2).
These consistent and detailed trends are observed not only over a 30 year period but also in evaluations over shorter periods such as 5 years. For example, in the IPC classes corresponding to Food chemistry, the trend of high TCI is consistent in the 2000s. Technologies with high scores as C13 and A21 and those that decline as A23 and A01 are captured in a consistent manner (SI Fig. A2).
Moreover, the corporate approach facilitates relative technological evaluations within specific regions, a capability absent in the traditional regional approach, which relies solely on the national-level presence of technologies. The corporate method identifies strategically important technologies within individual regions. For example, comparative analyses of TCI values for IPC classes among three representative prefectures, Tokyo, Osaka, and Aichi, highlight distinct local technological strengths. Aichi, notable for its concentration of major manufacturing corporations including Toyota, exhibits a technological profile markedly different from Tokyo and Osaka (Fig. 3). Technologies showing relatively higher local complexity despite lower overall national rankings underline region-specific corporate advantages.
@@
Figure 2: (A) The TCI scores of IPC first Class (N = 124) and Schmoch (N = 35) for both prefectures and corporations. The scores have been scaled to values ranging from 0 to 100. Each IPC is plotted against the corresponding Schmoch axis, and in cases of multiple correspondences (i.e., Schmoch category “Chemical Engineering” maps to IPC classes B01–B08, C14, etc.), each pair is plotted individually. (B) The distribution of the absolute differences between TCI of IPC and of Schmoch systems.
@@
Figure 3: Three scatter plots arranged in a single row compare the Corporate TCI in Japan (vertical axis) to the Corporate TCI in Tokyo (left panel), Osaka (middle panel), and Aichi (right panel) on the horizontal axis. Each axis is scaled from 0 to 100, and the black points represent data points for each classification or entity. The red dashed line indicates the diagonal (y = x).

### v5

# Measuring Technological Complexity Using Japanese Patents

@@ Abstract
As international competition intensifies in technologies, nations need to identify key technologies to foster innovation. However, the identification is challenging due to the independent and inherently complex nature of technologies. Traditionally, analyses of technological portfolios have been limited to simple evaluations, indicating merely whether a technology is specialized. We propose evaluating TCI at the corporate level, which provides finer granularity and more detailed insights than conventional regional evaluations by using Japanese patent data spanning fiscal years 1981 to 2010. Specifically, we analyze a bipartite network composed of 1,939 corporations connected to technological fields categorized into either 35 or 124 classifications. Our findings quantitatively characterize the ubiquity and sophistication of each technological field, reveal detailed technological trends reflecting broader societal contexts, and demonstrate methodological stability even when employing finer technological classifications. Additionally, our corporate-level analysis allows consistent comparisons across different regions and technological fields, clarifying regional advantages in specific technologies. The corporate level analysis also reveals a new possibility for formulating an innovation strategy.
@@
Technological Complexity, Patent Data, Bipartite Network
@@ Introduction
In the twenty-first century, intensifying international competition has made technological capability a central element of national competitiveness. Nations seek to foster environments that support the development of sophisticated technologies, as these capabilities are fundamental to sustainable economic growth. [EC_SSP_nodate, Teixeira2022, NISTEP_nodate] The ecosystems that produce such capabilities are, however, inherently complex, involving interactions among countries, corporations, and technologies [Weaver1948science]. This complexity makes it difficult to accurately identify the key technologies and actors within the system. Understanding this structure is crucial for developing effective science and technology strategies to strengthen a nation’s competitive advantage.
@@
Historically, the assessment of technological capabilities has predominantly relied on quantitative metrics derived from bibliographic data of patents and publications, ranging from simple patent counts to more complex citation-based indicators, including Technology Relevance (TR) [jaffe2019patent, aristodemou2018citations]. While such indicators offer certain insights by considering citation practices, patent age, and technological fields, they suffer from significant shortcomings. Their inherent dependence on the long-term accumulation of citations makes them susceptible to biases from citation behaviors, patent examiners, and strategic patenting. Crucially, these metrics fail to capture the core aspect of technological complexity, namely the intricate interplay among corporations and technologies. Consequently, approaches that rely solely on these data provide a limited perspective and inadequately reflect the qualitative dimensions of technological capability.
@@
The Hidalgo-Hausmann algorithm[hidalgo2009eci] provides a powerful solution by considering both the diversity of a region and the ubiquity of its products or technologies. This dual perspective allows the methodology to quantify their complexity, revealing detailed economic structures not captured by traditional metrics. The utility of this foundational framework has led to its extension into several distinct streams of research[Hidalgo2021, Hidalgo2023, Hausmann2024, Balland2022, Chakraborty2020]. One major development is the introduction of complementary metrics such as fitness-based approaches [Tacchella2012, Wu2016, Albeaik2017]. Another significant stream, which is the focus of this study, involves the adaptation of the algorithm to patent data, pioneered by Balland and Rigby to measure technological complexity [balland2017tci]. Building on this work, recent studies have further extended the patent-based analysis not only to Europe[Pinheiro2022, ballandandboschma2021], but also to Asian economies [Jun2023, Abay2024]. Despite this wide range of applications, the focus has predominantly remained at the regional level [Hidalgo2021]. Therefore, shifting the analytical focus to the corporate level, holds the potential to unlock finer-grained comparisons of technological capabilities and regional specializations, revealing dynamics that a regional-level view cannot capture.
@@
Our study applies the HH algorithm to an extensive set of Japanese patent records to quantify technological complexity at the corporation level. By constructing bipartite networks linking corporations to technology classes, we develop a detailed measure of technological sophistication. Our analysis investigates the relationships between technological complexity, ubiquity, and the average corporate diversity of patent technologies, providing new insights into the nature of advanced capabilities. The model developed at the corporation level provides a more robust and consistent assessment of technology than regional-level analyses, particularly for finely classified technological fields. Importantly, this study shows that a corporation-level approach enables a direct comparison of technological characteristics across different regions. By focusing on corporations within specific regions, we can identify distinct industrial specializations and reveal which technologies are complex within each region’s unique context. The analyses from this study provide insights for creating tailored industrial policies and more accurately identifying the key technologies vital for future growth.
@@
The rest of this paper is organized as follows. In Introduction section, we present the ubiquity and complexity of technological fields presented by the patent data, followed by a detailed analysis of technological trends. In the Consistency over Time section, we provide a critical analysis of our findings and discuss the implications of our results. Finally, in the Discussion section, we detail the data processing and methods used in our analysis.
@@ Results
In this study, we use the HH algorithm to quantify the structural characteristics of technologies based on corporate patent publications from 1,939 corporations filed between fiscal years 1981 and 2010 (see Discussion section). First, we define ubiquity (or degree centrality) $K_{T,0}$ as the number of corporations connected to a technology which lower value indicates that the technology is more rare. Similarly, the average diversity $K_{T,1}$, which measures the variety of corporations connected to a technology, implies that a higher value reflects diversified publication, whereas a lower value suggests specialization. Generally, the average of each indicator is used as a threshold to distinguish high from low values [hidalgo2009eci, balland2017tci, Balland2018].
@@
To systematically evaluate technological characteristics, we map the relationships among ubiquity $K_{T,0}$, average diversity $K_{T,1}$, and TCI across patent technologies (Fig. 1). Technologies are classified according to the Schmoch system [Schmoch2008], as illustrated in Panel A of Fig. 1. Values for ubiquity and average diversity are normalized by their respective means, while TCI is segmented by the zero threshold.
Figure 1: (A) The legend of the five classifications defined by Schmoch [Schmoch2008], along with an explanation of the technology codes used in Panels (B), (C), and (D). (B) The Ubiquity $K_{T,0}$ versus average diversity $K_{T,1}$. (C) The Ubiquity $K_{T,0}$ versus TCI. (D) The TCI versus average diversity $K_{T,1}$. The black lines in Panels (B), (C), and (D) indicate the mean values of $K_{T,0}$ and $K_{T,1}$, and $\text{TCI}=0$ is also shown with a black line. Each dot is color‐coded according to (A).
@@
Detailed analysis using International Patent Classification (IPC) classes further supports these observations (Fig. S1). Within the five major sectors defined by Schmoch [Schmoch2008], the relationship between ubiquity and average diversity is predominantly weak or moderate. Specifically, Electrical Engineering displays a strong positive correlation between ubiquity and TCI (Pearson correlation coefficient $r=0.738$), highlighting the coexistence of corporations specializing in broadly applicable technologies and those focusing on niche, specialized innovations.
@@
Electrical engineering technologies are predominantly located in the region of low TCI and low average diversity (Panel B in Fig. 1). In contrast, Chemistry and Pharmaceuticals technologies are clustered in the area of high TCI and high average diversity, exhibiting a positive correlation between these two metrics (Panel D in Fig. 1; Pearson correlation coefficient rr=0.316). Furthermore, corporations in the Chemistry and Pharmaceuticals domains are connected not only to fields with low ubiquity but also to those that are more common such as Biotechnology, Analysis of Biological Materials, and Measurement (Panel C in Fig. 1). Thus, these results demonstrate a fundamental difference between the two technological domains. The Chemistry and Pharmaceuticals sectors are characterized by rarity and high sophistication and are typically developed by more diversified corporations. A detailed examination reveals a slight negative correlation between ubiquity and average diversity (Fig. S1), suggesting that in sectors with high technological complexity, corporations with more comprehensive technological portfolios tend to produce rare technologies. Furthermore, the IPC classes corresponding to Biotechnology and Measurement show high TCI values in fields related to the production of sugar and yeast (C13, A21, C12), drug discovery (A61, C07), and the preservation of these technologies (C07, A23). These findings are consistent with the socio-technical context of the period, particularly reflecting the emergence of drug discovery technologies in the 1980s and 1990s [Sakakibara2014, Nakamura2022]. Moreover, the significant roles of food chemistry and biotechnology during the 1990s and 2000s underscore their contribution to the development of additives that enabled safer food preservation and processing not only in Japan but globally [Bhatia2018, Murakami2024].
Consistency over Time
The preceding section focuses on the corporate TCI over the full period from fiscal year 1981 to 2010 and provides a baseline view of technological characteristics in Japan. In this section, we explore whether more granular insights into regional technological capabilities can be obtained by comparing the regional- and corporate-level TCIs. Specifically, through comparing results derived from the Schmoch classes and the three-digit IPC classes, we examine how classification granularity affects the consistency of technological complexity evaluations across regions and corporations.
The two classification systems share the same technological scope but differ in their level of aggregation. The Schmoch system includes 35 broad technology classes, whereas the three digit IPC used in this study contains 124 finer classes. Previous studies on regional technological complexity have reported that using finer classifications tends to make the bipartite network sparse and to reduce the consistency of the calculated indicators[morrison2017economic, Tacchella2012]. Such inconsistency partly reflects the limited number of regional nodes relative to the number of technology classes. In regional analyses, several dozen to a few hundred regions such as countries or EU cities[Hidalgo2021, nepelski2020technological] are connected to several hundred or thousand technology classes, which limits the variation in ubiquity across technologies and produces a biased distribution of TCI. By contrast, corporate analyses involve a much larger number of nodes, often exceeding several thousand corporation, and therefore capture a broader range of ubiquity through the iterative calculation.
These tendencies are reflected in the observed distributions of TCI. In the regional analysis based on the Schmoch system, 26 out of 35 classes fall within the range of 75 to 100 (Fig. 2), indicating that most technologies exhibit similar levels of complexity. This limited dispersion is consistent with the small number of regional nodes, which comprises only 47 prefectures in Japan. In contrast, the corporate analysis involves 1,939 corporations, resulting in a denser bipartite network that captures greater heterogeneity in technological capabilities. Consequently, only 15 classes lie within the same range in the corporate case, while the remaining classes show a wider spread of TCI values across technologies. This broader dispersion suggests that the corporate analysis captures the variation in ubiquity more effectively, revealing distinct differences in how frequently each technology appears across corporations.
@@
Figure 2: (A) The TCI scores of 124 IPC the first three-digit classes and 35 Schmoch classes for both prefectures and corporations. The scores have been scaled to values ranging from 0 to 100. Each IPC is plotted against the corresponding Schmoch axis, and in cases of multiple correspondences (i.e., Schmoch class “Chemical Engineering” maps to IPC classes B01–B08, C14, etc.), IPC classes are plotted individually. (B) The distribution of the absolute differences between TCI of IPC and of Schmoch classes.
The difference between the approaches becomes even clearer when using a finer classification. When the classification becomes more detailed and the number of technology nodes increases, the regional approach cannot define ubiquity sufficiently. In the case of Civil engineering, the variance of TCI values for IPC classes corresponding to the same Schmoch class is larger than that for the corporate approach. The larger difference between the TCI values of the coarse classification and those of the finer classification indicates a greater inconsistency in the evaluation for the regional approach. A Mann-Whitney U test shows that the distribution of residuals between the TCI values of the Schmoch and IPC classes is significantly different (U = 7495.5, $p$ value = 6.55e-15, $\gamma$ value = 0.60), and that the residuals for the corporate analysis are smaller than those for the regional analysis (Panel B in Fig. 2). Smaller residuals imply that the corporate TCI is less affected by the choice of classification level and provides a more consistent measure of technological complexity. This result suggests that the corporate network captures the co-occurrence patterns of technologies more precisely because of the larger number of corporation nodes and the finer granularity of technological connections. Hence, the corporate analysis offers a more stable representation of how diverse and rare capabilities are combined in actual technological activities. These consistent and detailed trends are observed not only over the 30-year period but also in shorter five-year intervals (SI Fig. S2). For example, in the IPC classes related to Food Chemistry, a persistently high TCI is observed for sugars (C13) and baking technologies (A21) throughout the 2000s. By contrast, foodstuffs (A23) and agriculture-related technologies (A01) show gradual declines in ranking, particularly after the mid-1990s. This indicates that even within related technological fields, some areas maintain their complexity advantage while others lose ground. These results confirm that the relative positioning of technologies can be captured in a stable and systematic manner across different time resolutions.
Corporate analyses are useful not only for examining technological structures at finer classification and temporal scales but also for capturing spatial characteristics in greater detail than regional analyses. While previous regional studies have mainly compared aggregate indicators such as the average TCI or the presence of links to certain technologies, the corporate analysis enables a direct comparison of each technology’s TCI between different regions. Regional comparisons are informative because similarity or divergence in corporate TCI across prefectures can reveal distinctive industrial bases and specialization patterns. We compare cross-prefectural alignment by pairing Tokyo, Osaka, and Aichi, and interpret the class-level correspondence from the concentration around the diagonal and notable deviations in the scatter plots (Fig. 3). Tokyo and Osaka exhibit a close correspondence across many fields, which reflects a strong resemblance in their technological structures. Tokyo and Aichi show a looser association in which Aichi tends to be stronger in electronics and ICT, for example Computing (G06) and Telecommunications (H04), while Tokyo shows relative strengths in consumer goods and light manufacturing, for example Printing and Bookbinding (B42) and Lighting Technology (F21). The pair of Osaka and Aichi diverges the most, with Aichi again relatively strong in electronics, for example Computing (G06) and Basic Electronic Circuitry (H03), whereas Osaka appears higher in manufacturing-oriented areas, for example Presses (B30) and Decorative Arts and Processes (B44). This pattern may highlight that Aichi’s technological profile diverges not only from Tokyo but also from Osaka, reflecting a distinctive regional specialization in electronics and ICT. The corporate analysis thus makes finer regional technological characteristics visible within each prefecture than the regional-level analysis.
@@
Figure 3: Three scatter plots arranged in a single row compare the Corporate TCI between pairs of prefectures. The left panel compares Osaka with Tokyo, the middle panel compares Aichi with Tokyo, and the right panel compares Aichi with Osaka. Each axis ranges from 0 to 100. Black points represent individual classifications or entities, and the red dashed line indicates the diagonal reference line.